# Shapiro Delay Measurements from Fifteen Years of PSR J12311411 Radio Observations

### v1

@@ 4.2 PINT Gridding
The first set of analyses comprise PINT-based grids over various combinations of Shapiro delay parameters using all available GBT and Nançay data. This gridding routine is nearly identical to fon16 fon16, which was based on spl02 spl02; however, the fon16 implementation uses Tempo2. The version used in this work performs a minimization using PINT for given combinations of companion mass and inclination angle on a 200200 grid, holding cos() and fixed and fitting for all other timing model parameters. As in fon16 fon16, the map is transformed into a two-dimensional PDF and integrated along both axes to obtain one-dimensional PDFs for both Shapiro delay parameters.
@@
Figure 2 shows the resulting map and projected probability density functions for a trial that includes the white noise parameters in Table 1 (held fixed), parallax (PX), proper motion in RA and DEC (PMRA, PMDEC), an isotropic solar wind model that falls as (SWM0 as implemented in PINT), and as well as a Taylor-expansion model of the dispersion measure (DM) to two terms (DM1, DM2). With a minimum elongation of 9 degrees, delays from propagation in the solar wind are negligible. Employing alternate DM models (e.g. a single DM parameter or epoch-dependent DMX parameters) did not have an appreciable impact on the measured or cos() as seen in fon21; this is expected given the breadth of observing frequencies and high cadence of that data set. The inclusion of power-law red noise is not favored when these timing model parameters are present. From this trial, we find modest constraints on companion mass, inclination angle, and pulsar mass, with = M⊙, inclination = °, and = M⊙ (68.3% CI). Due to its faster speed, this method was effective for more thoroughly exploring alternate timing and noise models. Although it is possible to apply a prior to parameters in the gridding scheme, covariances between, e.g. binary and astrometric parameters, as well as the poorly-constrained companion mass, make the fully Bayesian approach described in Section 4.3 a more suitable choice for PSR J12311411.
4.3 PINT MCMC trials
@@
Bayesian approaches to pulsar timing offer numerous benefits over traditional pulsar timing techniques for a variety of specialized applications (e.g. vig14); however, the associated computational demand has prevented these methods from seeing widespread use. For example, the ability to include astrophysically motivated priors is a powerful improvement over typical least-squares techniques. This is of particular interest to astronomers wishing to conduct pulsar timing fits using complementary data sets (for example, one could use astrometric measurements from Gaia to constrain a pulsar’s parallax or proper motion). The PINT software package employs emcee, a widely-used open-source Python implementation of an affine-invariant ensemble sampler for MCMC (for13), to facilitate Bayesian timing model fits.
4.3.1 Priors on and sin()
@@
The evolutionary end-point of a low-mass white dwarf binary (that is, the system’s total mass and orbital period) can be directly predicted from the progenitor red giant star’s degenerate core mass and radius, and is strongly influenced by the star’s metallicity (ref71). In their 1999 work, Tauris & Savonije formulated a relationship between the binary orbital period of a pulsar system and the white dwarf companion’s mass (TS99; tau99 tau99, see also ist14 ist14 and ist16):
@@
Values of , , and are dependent on the chemical composition of the companion: Pop. I stars (young and metal-rich; = 0.02) have ( = 4.50, = , = 0.120); the median Pop. I/II (medium-) values are (4.75, , 0.115); and the Pop. II (old, metal-poor; = 0.001) values are (5.00, , 0.110). This relationship holds for companions with masses between 0.18 and 0.45 M⊙. A recent effort (mat20) extrapolates this relationship to lighter white dwarfs using the relationships outlined in ist16. For J12311411’s 1.86-day orbit, the TS99 Pop. I, Pop. I/II, and Pop. II values are 0.205, 0.214, and 0.223 M⊙, respectively. Priors on were chosen based on TS99 predictions (see Section 4.3.2).
@@
The data are informative themselves, and do not “require” the or sin() priors for a significant measurement. However, a flat prior on permits large, unphysical swaths of parameter space, and omitting the prior on sin() introduces an inaccurate bias in expected orbital orientations. In placing only modest priors on , we have elected to take a more conservative approach in an effort to understand the underlying physical processes rather than to obtain an improved pulsar mass constraint. However, in a context such as the X-ray lightcurve analysis of sal24, a restriction of parameter space to physical values is warranted (see Section 5).
4.3.2 Results of the Fixed-Noise MCMC Fits
We present the results from several MCMC trials using the radio data set. Each trial used 256 walkers and 3000 steps, with convergence ensured by calculating the autocorrelation time () for each chain in the fit (see goo10). Suggestions for chain lengths vary, though emcee recommends values . For each run, we have ensured that our integrated autocorrelation times and chain lengths meet that criterion, though many runs have factors of 1000 or more.
Figure 3 shows PDFs for , sin(), and the derived for two runs. The first (in orange), referred to as “TS99,” employs a wide Gaussian prior ( = 0.25 M⊙, bounded between 0 and 1 M⊙) on centered around the predicted intermediate-metallicity (Pop. I/II) companion mass value of 0.214 M⊙ as well as the “flat in cos()” prior. The second, designated “control,” uses flat priors for both parameters—a deliberately poor choice in the case of sin(). In assessing the prior, we also checked the impact of constraining the companion’s mass to a more physical maximum of 0.5 M⊙; however, the impact on is minimal () and the result is not plotted. The TS99 fit yields = M⊙, inclination = °, and = M⊙ (68.3% CI). The “control” (flat-prior) fit results in slightly worse constraints, with = M⊙, inclination = °, and = M⊙.
4.4 Results of the Fully-Bayesian Nonlinear MCMC Fits
@@
As a system with measurable but poorly constrained Shapiro delay, PSR J12311411 presented an opportunity to ensure no significant covariances between white noise parameters and timing model parameters exist, as well as glimpse what improvements might be possible with more sophisticated analysis. As in the PINT MCMC trials, we used the data set with the same JUMPs and timing model parameters. Pulsar mass values are constrained to a maximum of 3 M⊙, a significant difference from our previous trials. A prior on parallax determined by the pulsar’s DM distance (as in vig14 vig14, using the NE2001 electron density distribution from cor02 cor02) is applied, taking the form:
@@
As before, we find no significant red noise. Posterior distributions for all noise parameters measured in the a priori linear analysis and the generalized nonlinear analysis closely agree, with all median EQUAD and EFAC values varying by less than 5%. From this exploratory analysis, we find the companion mass = M⊙, inclination = °, and = M⊙ (68.3% CI). The derived mass and inclination values from the nonlinear runs agree with those from our MCMC trials, which is expected given the large uncertainties. However, we find that the new approach yields improved constraints on the orbital inclination and companion mass (see Figure 4). An observed decrease in is largely attributable to the imposition of a 3 M⊙ pulsar mass constraint. Taken together, these results suggest that in the case of PSR J12311411, determining the white and red noise values prior to MCMC fits is a reasonable approach (see Section 5). The improvements brought by this nonlinear technique merit further exploration, though the associated computational demand is high (see the upcoming paper by Kaiser et al.).
5 Discussion
@@
We explored a large number of model permutations during our analysis, such as the inclusion of poorly measured astrometric parameters, different approaches to DM and solar wind timeseries modeling, and different Shapiro delay parameterizations. We applied each of these in turn to the whole data set as well as its constituent backend-receiver-telescope members as individual data sets, though only results using the full data set are presented here (with the exception of Figures 6 and 7). While a detailed description of each of these trials is beyond the scope of this work, we found many to have a significant impact on measured white and intrinsic red noise, and in turn the pulsar and companion masses. We note that J12311411 is a somewhat unusual case in the landscape of precision pulsar mass measurements. In the “gold-standard” examples of Shapiro delay-derived masses, the systems (many of which are observed with pulsar timing arrays, and are therefore selected against strong scintillation or orbital variability) are sufficiently edge-on that the Shapiro delay effect is significantly more pronounced, rendering the mass measurement more independent of the analysis methods and assumed model.
@@
Because of PSR J12311411’s proximity and large proper motion, measurement of long-term secular orbital effects such as (change in binary orbital period), (change in projected semi-major axis), and (change in longitude of periastron) should be considered. For the first time to our knowledge, we report a significant and parallax measurement for J12311411. Observed is not solely (or even notably, in the case of PSR J12311411) attributable to a loss of orbital energy through gravitational radiation. Galactic differential rotation, z-axis acceleration within the Galactic potential, and the Shklovskii effect—a seeming acceleration due to astrometric parallax—all contribute to the total observed orbital period decay. All four of these effects are fundamentally defined by the individual component masses and distance to the system in addition to the well-measured timing parameters; therefore, an independent measurement of and with two significantly detected post-Keplerian parameters can be compared for consistency with a significantly measured parallax, as we see with PSR J12311411. Section 4.2 of fon21 fon21 (and references therein) presents a thorough overview of the methods and quantities used here, including explicit definitions for each component of . We have chosen to use the same Galactic speed and distance parameters cited in that work (gra19) as well as the traditional model of Galactic gravitational potential (kui89). We similarly do not include any negligible higher-order corrections to (hu20).
@@
Figure 5 summarizes the distance to PSR J12311411 as derived from radio measurements of and PX in addition to other relevant timing parameters. The DM-derived distance of 0.463 kpc from the NE2001 electron density model (as implemented in NE2001p by ock24 ock24) is also plotted for reference. Though is significantly better constrained than parallax, the two derived distance values are in agreement. At the time of the sal24 analysis, our radio timing model had erroneously excluded , as a recent addition of data caused the parameter to gain significance. Although the PX-derived distance constraint of 600100 pc was provided, a more conservative (uniform) prior between 100 and 700 pc was applied in the X-ray lightcurve modeling.
@@
As PSR J12311411’s orbit is nearly circular, secular variations in , , or due to intrinsic changes are not expected. However, just as an apparent was significantly measured, so too might the related non-relativistic (apparent) and . As in ala21, an F-test with p-value threshold of 0.0027 was conducted to determine whether the inclusion of these post-Keplerian parameters was appropriate. While was strongly inconsistent with noise, is borderline (2) and is not at all significant; therefore, the latter two parameters were not included in the model. The maximum contribution to from proper motion is given by
@@
where (the sum of the squares of proper motion in RA and DEC; see lor04 lor04). Our best-fit values of and yield a maximum due to proper motion of 2.04. As a point of reference, the current constraint on predicts a maximum inclination of (1 upper limit). With additional data, especially from next-generation facilities where timing precision will improve, may be significantly constrained.
@@
Continued Nançay timing may improve constraints on ; however, returns diminish as orbital phase becomes well-sampled. As an illustration, between MJD 56100 (the point at which Shapiro delay becomes measurable) and 56750—the first 1.8 years of observations—uncertainty in decreased by 1 M⊙. Between MJD 59000 and the data set’s last observation on 60408 (the last 3.8 years of data), uncertainty decreased by only 0.15 M⊙. Since MJD 59750, the most recent 1.8 years of data, uncertainty on is unchanged. An additional avenue for future exploration may be the use of PSR J12311411’s strong scintillation to obtain a series of scintillation arc measurements, which could yield constraints on the transverse velocity and in turn, independently measure the system’s inclination and longitude of ascending node (cor98).
@@
We find that the simultaneous determination of noise and timing model parameters described in Section 4.4 is better able to explore the parameter space when is not included and EQUAD is therefore high, yielding quite similar constraints on the Shapiro delay parameters regardless of the inclusion of . When noise is instead determined a priori and fixed during the MCMC sampling, the same flexibility is not observed; presumably because all MCMC samples have the same high EQUAD, the constraint is notably worse in the fits without .
@@
The measured PSR J12311411 companion mass = M⊙ is consistent with the TS99 prediction both with and without TS99 priors applied, but is not well-constrained enough to glimpse the system’s detailed evolutionary history. This long-term timing effort has indicated that, with a mildly-inclined orbit of °and inferred = M⊙ (or M⊙ from the simultaneous-sampling trials presented in Section 4.4), PSR J12311411 is much the same as typical MSP-white dwarf binaries.
However, these results—especially the relatively strong inclination constraint—still informed the NICER X-ray profile analysis (sal24). The mass-inclination priors employed in that work were based on a preliminary result using tighter priors in our MCMC fit (yielding = and inclination = °), and further modified by imposing a strict 1–3 M⊙ limit on the distribution, resampling, and applying further physical priors based on neutron star compactness. The resulting prior was essentially uniform over the chosen mass range. Given the faint Shapiro delay and troublesome radio timing for PSR J12311411, a broad mass prior is expected. However, our 5% constraint on orbital inclination, which is significantly more resilient to changes in fitting method, red or white noise inclusion, and chosen priors, could be considered the more informative addition to the analysis.
Posterior distributions for , , sin(), and PX for the MCMC trials presented in Section 4.3 and Section 4.3.2 are available on Zenodo: 10.5281/zenodo.17596801.
6 Conclusions

### v2

@@ IV.2 PINT Gridding
The first set of analyses comprise PINT-based grids over various combinations of Shapiro delay parameters using all available GBT and Nançay data. This gridding routine is nearly identical to Fonseca et al. 2016, which was based on Splaver et al. 2002; however, the Fonseca et al. (2016) implementation uses Tempo2. The version used in this work performs a minimization using PINT for given combinations of companion mass and inclination angle on a 200200 grid, holding cos() and fixed and fitting for all other timing model parameters. As in Fonseca et al. 2016, the map is transformed into a two-dimensional PDF and integrated along both axes to obtain one-dimensional PDFs for both Shapiro delay parameters.
@@
Figure 2 shows the resulting map and projected probability density functions for a trial that includes the white noise parameters in Table 2 (held fixed), parallax (PX), proper motion in RA and DEC (PMRA, PMDEC), an isotropic solar wind model that falls as (SWM0 as implemented in PINT), and as well as a Taylor-expansion model of the dispersion measure (DM) to two terms (DM1, DM2). With a minimum elongation of 9 degrees, delays from propagation in the solar wind are negligible. Employing alternate DM models (e.g. a single DM parameter or epoch-dependent DMX parameters) did not have an appreciable impact on the measured or cos() as seen in Fonseca et al. (2021); this is expected given the breadth of observing frequencies and high cadence of that data set. The inclusion of power-law red noise is not favored when these timing model parameters are present. From this trial, we find modest constraints on companion mass, inclination angle, and pulsar mass, with = M⊙, inclination = °, and = M⊙ (68.3% CI). Due to its faster speed, this method was effective for more thoroughly exploring alternate timing and noise models. Although it is possible to apply a prior to parameters in the gridding scheme, covariances between, e.g. binary and astrometric parameters, as well as the poorly-constrained companion mass, make the fully Bayesian approach described in Section IV.3 a more suitable choice for PSR J12311411.
| Backend & Receiver | EFAC | EQUAD (s) | JUMP () |
| GBT 350 MHz | 1.360 | 0.053 | |
| GBT 820 MHz (2009) | 1.745 | 0.079 | |
| GBT 820 MHz (2019) | 1.146 | 3.838 | |
| NRT L-band, BON (2009) | 0.691 | 0.234 | |
| NRT L-band, NUPPI (pre-58600) | 1.061 | 0.073 | – |
| NRT L-band, NUPPI (post-58600) | 1.064 | 0.124 |
IV.3 PINT MCMC trials
@@
Bayesian approaches to pulsar timing offer numerous benefits over traditional pulsar timing techniques for a variety of specialized applications (e.g. Vigeland and Vallisneri, 2014); however, the associated computational demand has prevented these methods from seeing widespread use. For example, the ability to include astrophysically motivated priors is a powerful improvement over typical least-squares techniques. This is of particular interest to astronomers wishing to conduct pulsar timing fits using complementary data sets (for example, one could use astrometric measurements from Gaia to constrain a pulsar’s parallax or proper motion). The PINT software package employs emcee, a widely-used open-source Python implementation of an affine-invariant ensemble sampler for MCMC (Foreman-Mackey et al., 2013), to facilitate Bayesian timing model fits.
IV.3.1 Priors on and sin()
@@
The evolutionary end-point of a low-mass white dwarf binary (that is, the system’s total mass and orbital period) can be directly predicted from the progenitor red giant star’s degenerate core mass and radius, and is strongly influenced by the star’s metallicity (Refsdal and Weigert, 1971). In their 1999 work, Tauris & Savonije formulated a relationship between the binary orbital period of a pulsar system and the white dwarf companion’s mass (TS99; Tauris and Savonije 1999, see also Istrate et al. 2014 and 2016):
@@
Values of , , and are dependent on the chemical composition of the companion: Pop. I stars (young and metal-rich; = 0.02) have ( = 4.50, = , = 0.120); the median Pop. I/II (medium-) values are (4.75, , 0.115); and the Pop. II (old, metal-poor; = 0.001) values are (5.00, , 0.110). This relationship holds for companions with masses between 0.18 and 0.45 M⊙. A recent effort (Mata Sánchez et al., 2020) extrapolates this relationship to lighter white dwarfs using the relationships outlined in Istrate et al. (2016). For J12311411’s 1.86-day orbit, the TS99 Pop. I, Pop. I/II, and Pop. II values are 0.205, 0.214, and 0.223 M⊙, respectively. Priors on were chosen based on TS99 predictions (see Section IV.3.2).
@@
The data are informative themselves, and do not “require” the or sin() priors for a significant measurement. However, a flat prior on permits large, unphysical swaths of parameter space, and omitting the prior on sin() introduces an inaccurate bias in expected orbital orientations. In placing only modest priors on , we have elected to take a more conservative approach in an effort to understand the underlying physical processes rather than to obtain an improved pulsar mass constraint. However, in a context such as the X-ray lightcurve analysis of Salmi et al. (2024), a restriction of parameter space to physical values is warranted (see Section V).
IV.3.2 Results of the Fixed-Noise MCMC Fits
We present the results from several MCMC trials using the radio data set. Each trial used 256 walkers and 3000 steps, with convergence ensured by calculating the autocorrelation time () for each chain in the fit (see Goodman and Weare, 2010). Suggestions for chain lengths vary, though emcee recommends values . For each run, we have ensured that our integrated autocorrelation times and chain lengths meet that criterion, though many runs have factors of 1000 or more.
Figure 3 shows PDFs for , sin(), and the derived for two runs. The first (in orange), referred to as “TS99,” employs a wide Gaussian prior ( = 0.25 M⊙, bounded between 0 and 1 M⊙) on centered around the predicted intermediate-metallicity (Pop. I/II) companion mass value of 0.214 M⊙ as well as the “flat in cos()” prior. The second, designated “control,” uses flat priors for both parameters—a deliberately poor choice in the case of sin(). In assessing the prior, we also checked the impact of constraining the companion’s mass to a more physical maximum of 0.5 M⊙; however, the impact on is minimal () and the result is not plotted. The TS99 fit yields = M⊙, inclination = °, and = M⊙ (68.3% CI). Imposing an additional prior disallowing pulsar mass values greater than 3 M⊙ yields = M⊙, inclination = °, and = M⊙ (see Figure 4). The “control” (flat-prior) fit results in slightly worse constraints, with = M⊙, inclination = °, and = M⊙.
IV.4 Results of the Fully-Bayesian Nonlinear MCMC Fits
@@
As a system with measurable but poorly constrained Shapiro delay, PSR J12311411 presented an opportunity to ensure no significant covariances between white noise parameters and timing model parameters exist, as well as glimpse what improvements might be possible with more sophisticated analysis. As in the PINT MCMC trials, we used the data set with the same JUMPs and timing model parameters. Pulsar mass values are constrained to a maximum of 3 M⊙, a significant difference from our previous trials. A prior on parallax determined by the pulsar’s DM distance (as in Vigeland and Vallisneri 2014, using the NE2001 electron density distribution from Cordes and Lazio 2002) is applied, taking the form:
@@
As before, we find no significant red noise. Posterior distributions for all noise parameters measured in the a priori linear analysis and the generalized nonlinear analysis closely agree, with all median EQUAD and EFAC values varying by less than 5%. From this exploratory analysis, we find the companion mass = M⊙, inclination = °, and = M⊙ (68.3% CI). The derived mass and inclination values from the nonlinear runs agree with those from our MCMC trials, which is expected given the large uncertainties. However, we find that the new approach yields improved constraints on the orbital inclination and companion mass (see Figure 4). An observed decrease in is largely attributable to the imposition of a 3 M⊙ pulsar mass constraint, which is reflected in the close agreement seen between the fully generalized measurements and the fixed-noise MCMC results when this maximum mass prior is applied. Taken together, these results suggest that in the case of PSR J12311411, determining the white and red noise values prior to MCMC fits is a reasonable approach (see Section V). The improvements brought by this nonlinear technique merit further exploration, though the associated computational demand is high (see the upcoming paper by Kaiser et al.).
V Discussion
@@
We explored a large number of model permutations during our analysis, such as the inclusion of poorly measured astrometric parameters, different approaches to DM and solar wind timeseries modeling, and different Shapiro delay parameterizations. We applied each of these in turn to the whole data set as well as its constituent backend-receiver-telescope members as individual data sets, though only results using the full data set are presented here (with the exception of several illustrative comparisons presented later in this section). While a detailed description of each of these trials is beyond the scope of this work, we found many to have a significant impact on measured white and intrinsic red noise, and in turn the pulsar and companion masses. We note that J12311411 is a somewhat unusual case in the landscape of precision pulsar mass measurements. In the “gold-standard” examples of Shapiro delay-derived masses, the systems (many of which are observed with pulsar timing arrays, and are therefore selected against strong scintillation or orbital variability) are sufficiently edge-on that the Shapiro delay effect is significantly more pronounced, rendering the mass measurement more independent of the analysis methods and assumed model.
@@
Because of PSR J12311411’s proximity and large proper motion, measurement of long-term secular orbital effects such as (change in binary orbital period), (change in projected semi-major axis), and (change in longitude of periastron) should be considered. For the first time to our knowledge, we report a significant and parallax measurement for J12311411. Observed is not solely (or even notably, in the case of PSR J12311411) attributable to a loss of orbital energy through gravitational radiation. Galactic differential rotation, z-axis acceleration within the Galactic potential, and the Shklovskii effect—a seeming acceleration due to astrometric parallax—all contribute to the total observed orbital period decay. All four of these effects are fundamentally defined by the individual component masses and distance to the system in addition to the well-measured timing parameters; therefore, an independent measurement of and with two significantly detected post-Keplerian parameters can be compared for consistency with a significantly measured parallax, as we see with PSR J12311411. Section 4.2 of Fonseca et al. 2021 (and references therein) presents a thorough overview of the methods and quantities used here, including explicit definitions for each component of . We have chosen to use the same Galactic speed and distance parameters cited in that work (GRAVITY Collaboration et al., 2019) as well as the traditional model of Galactic gravitational potential (Kuijken and Gilmore, 1989). We similarly do not include any negligible higher-order corrections to (Hu et al., 2020).
@@
Figure 5 summarizes the distance to PSR J12311411 as derived from radio measurements of and PX in addition to other relevant timing parameters. The DM-derived distance of 0.463 kpc from the NE2001 electron density model (as implemented in NE2001p by Ocker and Cordes 2024) is also plotted for reference. Though is significantly better constrained than parallax, the two derived distance values are in agreement. At the time of the Salmi et al. (2024) analysis, our radio timing model had erroneously excluded , as a recent addition of data caused the parameter to gain significance. Although the PX-derived distance constraint of 600100 pc was provided, a more conservative (uniform) prior between 100 and 700 pc was applied in the X-ray lightcurve modeling.
@@
As PSR J12311411’s orbit is nearly circular, secular variations in , , or due to intrinsic changes are not expected. However, just as an apparent was significantly measured, so too might the related non-relativistic (apparent) and . As in Alam et al. (2021), an F-test with p-value threshold of 0.0027 was conducted to determine whether the inclusion of these post-Keplerian parameters was appropriate. While was strongly inconsistent with noise, is borderline (2) and is not at all significant; therefore, the latter two parameters were not included in the model. The maximum contribution to from proper motion is given by
@@
where (the sum of the squares of proper motion in RA and DEC; see Lorimer and Kramer 2004). Our best-fit values of and yield a maximum due to proper motion of 2.04. As a point of reference, the current constraint on predicts a maximum inclination of (1 upper limit). With additional data, especially from next-generation facilities where timing precision will improve, may be significantly constrained.
@@
Continued Nançay timing may improve constraints on ; however, returns diminish as orbital phase becomes well-sampled. As an illustration, between MJD 56100 (the point at which Shapiro delay becomes measurable) and 56750—the first 1.8 years of observations—uncertainty in decreased by 1 M⊙. Between MJD 59000 and the data set’s last observation on 60408 (the last 3.8 years of data), uncertainty decreased by only 0.15 M⊙. Since MJD 59750, the most recent 1.8 years of data, uncertainty on is unchanged. An additional avenue for future exploration may be the use of PSR J12311411’s strong scintillation to obtain a series of scintillation arc measurements, which could yield constraints on the transverse velocity and in turn, independently measure the system’s inclination and longitude of ascending node (Cordes and Rickett, 1998).
@@
We find that the simultaneous determination of noise and timing model parameters described in Section IV.4 is better able to explore the parameter space when is not included and EQUAD is therefore high, yielding quite similar constraints on the Shapiro delay parameters regardless of the inclusion of . When noise is instead determined a priori and fixed during the MCMC sampling, the same flexibility is not observed; presumably because all MCMC samples have the same high EQUAD, the constraint is notably worse in the fits without .
@@
The measured PSR J12311411 companion mass = M⊙ is consistent with the TS99 prediction both with and without TS99 priors applied, but is not well-constrained enough to glimpse the system’s detailed evolutionary history. Our findings are also in agreement with analyses of optical photometric and spectroscopic observations of PSR J12311411’s companion using the Gemini South telescope as presented in Testa et al. 2015 and Bassa et al. 2016. The former discovered the optical counterpart, finding that it must be cool and old ( K and age Gyr), with a distance kpc. Bassa et al. 2016 reported that companion masses between 0.188–0.384 M⊙ and distances 0.35–0.51 kpc are the allowable ranges to reproduce the optical and radio results using their chosen helium white dwarf models. If the companion is confirmed to have K, it could be one of the coolest white dwarfs known. The optical results, taken together with our long-term timing effort, have indicated that with a mildly-inclined orbit of °and inferred = M⊙ (or M⊙ from the simultaneous-sampling trials presented in Section IV.4), PSR J12311411 is likely a typical MSP-white dwarf binary.
Despite large uncertainties on and , these results—especially the relatively strong inclination constraint—still informed the NICER X-ray profile analysis (Salmi et al., 2024). The mass-inclination priors employed in that work were based on a preliminary result using tighter priors in our MCMC fit (yielding = and inclination = °), and further modified by imposing a strict 1–3 M⊙ limit on the distribution, resampling, and applying further physical priors based on neutron star compactness. The resulting prior was essentially uniform over the chosen mass range. Given the faint Shapiro delay and troublesome radio timing for PSR J12311411, a broad mass prior is expected. However, our 5% constraint on orbital inclination, which is significantly more resilient to changes in fitting method, red or white noise inclusion, and chosen priors, could be considered the more informative addition to the analysis.
Posterior distributions for , , sin(), and PX for the MCMC trials presented in Section IV.3 and Section IV.3.2 are available on Zenodo: 10.5281/zenodo.17596801.
VI Conclusions